# The Automorphism Equivariant Hitchin Index

### v1

@@ 2.3 Identification of The Torus Fixed Locus
$1$ The construction of the isomorphism
$\rho_{m}:X_{m}^{f}\rightarrow\bigsqcup_{j\in\mathcal{J}_{m}}\tilde{X}_{l_{j}},$ (61)
is straightforward, and we have $X^{f}_{m,j}=\rho_{m}^{-1}(\tilde{X}_{l_{j}})$ for all $j\in\mathcal{J}_{m}$. For $\tilde{D}\in\tilde{X}_{l_{j}}$, the inverse is given by $\tilde{D}\mapsto\pi^{*}(\tilde{D})+j.$ The proof that this is indeed an isomorphism uses that $p$ is a prime and is elementary. We will now show that $\delta:\mathcal{F}_{m}^{f}\rightarrow X_{m}^{f}$ is the projection of a principal $\operatorname{Jac}(\tilde{X})[2]$-bundle. The decomposition given in equation (61) will then entail that $\mathcal{M}_{i,j}^{\mathbb{T}}$, as defined in equation (59), is in fact a principal $\operatorname{Jac}(\tilde{X})[2]$-bundle over $\tilde{X}_{l_{j}}$. The proof of the equivariant decomposition given in equation (60) will be given in the third step of the proof, and this will imply that $\mathcal{M}_{i,j}^{\mathbb{T}}$ is a subspace of $\mathcal{M}_{i}$. First we notice that $\gcd(p,2)=1$ implies that $\pi^{*}(\operatorname{Jac}(X))[2]=\operatorname{Jac}(X)[2]^{f}$, and it is elementary to verify that $\pi^{*}$ defines an isomorphism. Now notice that the image of $\operatorname{\mathcal{F}}_{m}^{f}$ under $\delta:\operatorname{\mathcal{F}}_{m}\rightarrow X_{m}$ belong to $X_{m}^{f}$. Further, if $\delta(\mathcal{E}_{1})=\delta(\mathcal{E}_{2})$ for $\operatorname{\mathcal{E}}_{1},\operatorname{\mathcal{E}}_{2}\in\operatorname{\mathcal{F}}_{m}^{f}$, then there exists a unique $L\in\operatorname{Jac}(X)[2]$ with $L.\mathcal{E}_{1}=\mathcal{E}_{2}$. Then
@@
and therefore the fact that $\operatorname{\mathcal{F}}_{m}$ is a $\operatorname{Jac}(X)[2]$-torsor implies that $f^{*}(L)=L$. Thus $L\in\operatorname{Jac}(X)[2]^{f}$. Therefore the Lemma will follow if we can establish surjectivity of the map $\delta:\operatorname{\mathcal{F}}_{m}^{f}\rightarrow X_{m}^{f}$. Given $\delta_{0}\in X_{m}^{f}$ we want to find $L_{0}\in\operatorname{Pic}(X)^{f}$, that solves the equation
@@
Thus we have reduced the problem to finding an $\langle f\rangle$-equivariant square-root of $\Lambda(-\delta_{0})$. Using the first part of the lemma, we see that there exists $j\in\mathcal{J}_{m}$ and $D\in\tilde{X}_{l_{j}}$, such that $\delta_{0}=\pi^{*}(D)+j$. As $p$ is an odd prime, we have $2\in(\mathbb{Z}/p\mathbb{Z})^{\times}$, and therefore we can for each $x\in X^{f}$ find $a_{x}\in\mathbb{Z}$ such that
@@
$L=\mathcal{O}_{-(j+2A)}/\langle f\rangle\in\operatorname{Pic}(\tilde{X}),$ (65)
and we have that $\mathcal{O}_{-(j+2A)}=\pi^{*}(L)$. Recall that $\Lambda=\pi^{*}(\tilde{\Lambda})$. We observe that
@@
$2$. Let $x\in X^{f}$. As $f_{\operatorname{\mathbbm{E}}}$ commutes with the $\mathbb{T}$-action, and as $\mathbb{T}$ act with weight $t$ on $\operatorname{\mathbbm{E}}_{m,1}$ and with weight equal to unity on $\operatorname{\mathbbm{E}}_{m,0}$, it easily follows that we must have a $\mathbb{T}\times\langle f\rangle$-equivariant decomposition of the form (60), for some $\mu^{\prime}_{x}\in\mathbb{Z}/p\mathbb{Z}$ in place of $\epsilon_{j,x}i_{x}$. Thus we must prove that
@@
Let $\operatorname{\mathcal{E}}=(L_{1}\oplus L_{2},\Phi_{b})\in\mathcal{M}_{i,j}^{\mathbb{T}}$ be given. Consider a frame $(s_{1},s_{2})$ near $x$, where $s_{j}$ is a frame of $L_{j}$ for $j=1,2$. Let $z$ be a coordinate centered at $x$. Write $b=\phi(s_{2}\otimes s_{1}^{*})dz$, where $\phi\in\mathcal{O}_{X,x}$ is a regular function vanishing to order $d$ at $x$. Then $d=j_{x}\mod p$ by definition. Thus we may write $\phi=z^{d}\cdot h_{1}(z)$ with $h_{1}(x)\not=0.$ It follows that $f^{*}(\phi)=\zeta^{n_{x}j_{x}}z^{d}f^{*}(h_{1})$. As $df^{*}(f^{*}(dz))=\frac{\partial f}{\partial z}dz=\zeta^{n_{x}}dz$, we obtain
@@
As $\operatorname{\mathcal{E}}$ is a fixed point, equation (19) holds for $\Phi_{b}$, and thus the right hand side of equation (68) and the right hand side of equation (69) are equal, i.e. we have that
@@
Fix $i\in\mathcal{I}.$ Towards the end of identifying the components of ~$\tilde{\mathcal{M}}_{i}^{\mathbb{T}}$, we note that $\tilde{\mathcal{N}}_{i}$ is always one of the components, so it suffices to identify the components of ~$\tilde{\mathcal{M}}_{i}^{\mathbb{T}}\setminus\tilde{\mathcal{N}}_{i}$. By the above lemma this is equivalent to identifying the image $\varXi_{i}(\mathcal{M}_{i,j}^{\mathbb{T}})$ for all $m\in M$ and $j\in\mathcal{J}_{i,m}.$
@@
Set $\mathcal{J}^{\prime}_{i}=\sqcup_{m\in M}\mathcal{J}_{i,m}$. For $j\in\mathcal{J}^{\prime}_{i}$ denote by ~$\tilde{\mathcal{M}}_{i,j}^{\mathbb{T}}$ the set of isomorphism classes of quasi parabolic Higgs bundles $(E,F,\Phi_{b})$ on $(\tilde{X},\operatorname{\tilde{\mathcal{D}}}_{i})$ of rank two of the following type.
  • There are $L_{+},L_{-}\in\operatorname{Pic}(\tilde{X})$ with $E=L_{+}\oplus L_{-}$ and $L_{+}L_{-}\simeq\tilde{\Lambda}_{i}$, and for each $\tilde{x}\in\operatorname{\tilde{\mathcal{D}}}_{i}$, we have that $F(\tilde{x})=L_{\epsilon_{j,x}}(\tilde{x})$.
  • The Higgs field $\Phi_{b}$ is the off-diagonal Higgs field associated with a section $b\in H^{0}(\tilde{X},L_{+}L_{-}^{-1}K_{\tilde{X}}(\operatorname{\tilde{\mathcal{D}}}_{i}))$. The zero divisor of $b$ has degree $l_{j}$.
@@
Fix $i\in\mathcal{I}$. Set $\mathcal{J}_{i}=\{i\}\sqcup\mathcal{J}_{i}^{\prime}$ and set ~$\tilde{\mathcal{M}}^{\mathbb{T}}_{i,i}=\tilde{\mathcal{N}}_{i}$. Fix $j\in\mathcal{J}^{\prime}_{i}$. Then $\varXi(\mathcal{M}^{\mathbb{T}}_{i,j})=\tilde{\mathcal{M}}^{\mathbb{T}}_{i,j}$. Thus the set of components of ~$\tilde{\mathcal{M}}_{i}^{\mathbb{T}}$ is in bijection with $\mathcal{J}_{i}$. In particular, every ~$\operatorname{\mathcal{E}}\in\tilde{\mathcal{M}}_{i,j}^{\mathbb{T}}$ is stable with resepct to the weight $w_{i}$.
@@
The proof follows from an elementary analysis of how the Hecke modification $\Psi_{i}$ act on $\operatorname{\mathcal{E}}=(L_{1}\oplus L_{2},\Phi_{b})\in\mathcal{M}_{i,j}^{\mathbb{T}}$. Consider the setup of the second step of the proof of Proposition 2.4. Let $x\in\mathcal{D}_{i}$. Assume for definiteness that $\epsilon_{j,x}=1$. We can choose the frame $(s_{1},s_{2})$ of $\operatorname{\mathcal{E}}$ such that $s_{1}$ is a non-vanishing section of $L_{1}$ and $s_{2}$ is a non-vanishing section of $L_{2}$. Then the Hecke transform is given as follows
$\operatorname{\mathcal{E}}^{\prime}=(L_{1}\oplus L_{2}(-x),\Phi_{b}^{\prime}).$ (71)
Write $E=L_{1}\oplus L_{2}$ and $E^{\prime}=L_{1}\oplus L_{2}(-x)$. Notice that in the frame $(s_{1},s_{2})$, the matrix of $f_{E}$ is diagonal in a neighbourhood of $x$. Therefore equation (30) implies that the lift $f_{E^{\prime}}$ of the Hecke transform $\operatorname{\mathcal{E}}^{\prime}$ is diagonal with respect to any frame of the form $(s_{1}^{\prime},s_{2}^{\prime})$ where $s_{1}^{\prime}$ is a non-vanishing section of $L_{1}$ and $s_{2}^{\prime}$ is a non-vanishing section of $L_{2}(-x)$. Similarly, equation (27) implies that $\Phi^{\prime}_{b}$ will also be an off-diagonal Higgs field associated with a section.
@@
That every ~$\operatorname{\mathcal{E}}\in\tilde{\mathcal{M}}_{i,j}^{\mathbb{T}}$ is stable with respect to $w_{i}$ follows from Theorem 1.1, the fact that $\mathcal{M}^{\mathbb{T}}_{i,j}$ is a sub-variety of $\mathcal{M}_{i}$ and the identity $\varXi(\mathcal{M}^{\mathbb{T}}_{i,j})=\tilde{\mathcal{M}}^{\mathbb{T}}_{i,j}$. ∎
@@
From Corollary 2.1 it follows that every connected component is either a smooth moduli space of stable parabolic vector bundles on $\tilde{X}$ of rank two, or a principal-$\operatorname{Jac}(\tilde{X})[2]$ bundle over a symmetric power of $\tilde{X}$. Consider first the case of moduli spaces of stable bundles. The intersection paring of these moduli spaces have been thoroughly studied following Witten’s work [66]. In the so-called coprime case, formulae for all pairings were proven in rank two in [59] and in all ranks in [42]. Subsequently, a version of these formulae for a class of parabolic moduli spaces including $\tilde{\mathcal{N}}_{i}$ were proven in [meinrenkenWittenFormulaeIntersection2005]. Consider the case of the intersection pairings on the cohomology ring of symmetric powers of a Riemann surface. These rings are generated by MacDonald classes [49] and the intersection pairings were computed by Zagier, and the resulting formulae can be found in [60].

### v2

@@ 1.3 Identification of The Torus Fixed Locus
1.1. The construction of the isomorphism
$\rho_{c}:X_{c}^{f}\rightarrow\bigsqcup_{j\in\mathcal{J}_{c}}\tilde{X}_{l_{j}},$ (43)
is straightforward, and we have $X^{f}_{c,j}=\rho_{c}^{-1}(\tilde{X}_{l_{j}})$ for all $j\in\mathcal{J}_{c}$. For $\tilde{D}\in\tilde{X}_{l_{j}}$, the inverse is given by $\tilde{D}\mapsto\pi^{*}(\tilde{D})+j.$ The proof that this is indeed an isomorphism is elementary.
We will now show that $\delta:\operatorname{\mathcal{F}}_{c}^{f}\rightarrow X_{c}^{f}$ is the projection of a principal $\operatorname{Jac}(\tilde{X})[2]$-bundle. The decomposition given in equation (43) will then entail that $\mathcal{M}_{i,j}^{\mathbb{T}}$, as defined in equation (41), is in fact a principal $\operatorname{Jac}(\tilde{X})[2]$-bundle over $\tilde{X}_{l_{j}}$. The proof of the equivariant decomposition given in equation (42) will be given in the third step of the proof, and this will imply that $\mathcal{M}_{i,j}^{\mathbb{T}}$ is a subspace of $\mathcal{M}_{i}$.
First we notice that $\gcd(p,2)=1$ implies that $\pi^{*}(\operatorname{Jac}(X))[2]=\operatorname{Jac}(X)[2]^{f}$, and it is elementary to verify that $\pi^{*}$ defines an isomorphism. Now notice that the image of $\operatorname{\mathcal{F}}_{c}^{f}$ under $\delta:\operatorname{\mathcal{F}}_{c}\rightarrow X_{c}$ belong to $X_{c}^{f}$. Further, if $\delta(\mathcal{E}_{1})=\delta(\mathcal{E}_{2})$ for $\operatorname{\mathcal{E}}_{1},\operatorname{\mathcal{E}}_{2}\in\operatorname{\mathcal{F}}_{c}^{f}$, then there exists a unique $L\in\operatorname{Jac}(X)[2]$ with $L.\mathcal{E}_{1}=\mathcal{E}_{2}$. Then
@@ 1.1. The construction of the isomorphism
and therefore the fact that $\operatorname{\mathcal{F}}_{c}$ is a $\operatorname{Jac}(X)[2]$-torsor implies that $f^{*}(L)=L$. Thus $L\in\operatorname{Jac}(X)[2]^{f}$. Therefore the lemma will follow if we can establish surjectivity of the map $\delta:\operatorname{\mathcal{F}}_{c}^{f}\rightarrow X_{c}^{f}$. Given $\delta_{0}\in X_{c}^{f}$ we want to find $L_{0}\in\operatorname{Pic}(X)^{f}$, that solves the equation
@@
Thus we have reduced the problem to finding an $\langle f\rangle$-equivariant square-root of $\Lambda(-\delta_{0})$. Using the first part of the lemma, we see that there exists $j\in\mathcal{J}_{c}$ and $D\in\tilde{X}_{l_{j}}$, such that $\delta_{0}=\pi^{*}(D)+j$. As $p$ is an odd prime, we have $2\in(\mathbb{Z}/p\mathbb{Z})^{\times}$, and therefore we can for each $x\in X^{f}$ find $a_{x}\in\mathbb{Z}$ such that
@@
[…]
@@
22. Recall the content of Remark 2. Let $x\in X^{f}$. As $f_{\operatorname{\mathbbm{E}}}$ commutes with the $\mathbb{T}$-action, and as $\mathbb{T}$ act with weight $t$ on $\operatorname{\mathbbm{E}}_{c,1}$ and with weight equal to unity on $\operatorname{\mathbbm{E}}_{c,0}$, it easily follows that we must have a $\mathbb{T}\times\langle f\rangle$-equivariant decomposition of the form (42), for some $\mu^{\prime}_{x}\in\mathbb{Z}/p\mathbb{Z}$ in place of $\epsilon_{j,x}i_{x}$. Thus we must prove that
@@
Let $\operatorname{\mathcal{E}}=(L_{1}\oplus L_{2},\Phi_{b})\in\mathcal{M}_{i,j}^{\mathbb{T}}$ be given. Consider a frame $(s_{1},s_{2})$ near $x$, where $s_{j}$ is a frame of $L_{j}$ for $j=1,2$. Let $z$ be a coordinate centered at $x$. Write $b=\phi(s_{2}\otimes s_{1}^{*})dz$, where $\phi\in\mathcal{O}_{X,x}$ is a regular function vanishing to order $d$ at $x$. Then $d=j_{x}\mod p$ by definition. Thus we may write $\phi=z^{d}\cdot h_{1}(z)$ with $h_{1}(0)\not=0.$ It follows that $f^{*}(\phi)=\zeta^{n_{x}j_{x}}z^{d}f^{*}(h_{1})$. As $df^{*}(f^{*}(dz))=\frac{\partial f}{\partial z}dz=\zeta^{n_{x}}dz$, we obtain
@@
[…]
@@
Fix $i\in\mathcal{I}.$ Towards the end of identifying the components of $\tilde{\mathcal{M}}_{i}^{\mathbb{T}}$, we note that $\tilde{\mathcal{N}}_{i}$ is always one of the components, so it suffices to identify the components of $\tilde{\mathcal{M}}_{i}^{\mathbb{T}}\setminus\tilde{\mathcal{N}}_{i}$. By the above lemma this is equivalent to identifying the image $\varXi_{i}(\mathcal{M}_{i,j}^{\mathbb{T}})$ for all $c\in C$ and $j\in\mathcal{J}_{i,c}.$ For all $c\in C$ and $j\in\mathcal{J}_{i,c}$ define the effective divisor $\operatorname{\tilde{\mathcal{D}}}_{ij}$ on $\tilde{X}$ by
$\operatorname{\tilde{\mathcal{D}}}_{ij}:=\sum_{x\in\mathcal{D}_{i}:\epsilon_{j,x}=-1}\pi(x).$ (53)
Below, we implicitly use the fact that the sheaf of holomorphic sections of $K_{\tilde{X}}(\operatorname{\tilde{\mathcal{D}}}_{i}-\operatorname{\tilde{\mathcal{D}}}_{ij})$ is a subsheaf of the sheaf of holomorphic sections of $K_{\tilde{X}}(\operatorname{\tilde{\mathcal{D}}}_{i})$. This fact implies that for any holomorphic line bundle $L\rightarrow\tilde{X}$, there is a natural inclusion
$H^{0}(\tilde{X},LK_{\tilde{X}}(\operatorname{\tilde{\mathcal{D}}}_{i}-\operatorname{\tilde{\mathcal{D}}}_{ij}))\hookrightarrow H^{0}(\tilde{X},LK_{\tilde{X}}(\operatorname{\tilde{\mathcal{D}}}_{i})).$ (54)
@@
Set $\mathcal{J}^{\prime}_{i}=\sqcup_{c\in C}\mathcal{J}_{i,c}$. For $j\in\mathcal{J}^{\prime}_{i}$ denote by $\tilde{\mathcal{M}}_{i,j}^{\mathbb{T}}$ the set of isomorphism classes of quasi parabolic Higgs bundles $(E,F,\Phi_{b})$ on $(\tilde{X},\operatorname{\tilde{\mathcal{D}}}_{i})$ of rank two of the following type.
[…]
  • The Higgs field $\Phi_{b}$ is the off-diagonal Higgs field associated with a section $b\in H^{0}(\tilde{X},L_{+}L_{-}^{-1}K_{\tilde{X}}(\operatorname{\tilde{\mathcal{D}}}_{i}-\operatorname{\tilde{\mathcal{D}}}_{ij}))$. The zero divisor of $b$ has degree $l_{j}$.
Remark 4.
[…]
@@
Fix $i\in\mathcal{I}$. Set $\tilde{\mathcal{M}}^{\mathbb{T}}_{i,i}=\tilde{\mathcal{N}}_{i}$. Fix $j\in\mathcal{J}^{\prime}_{i}$. Then $\varXi(\mathcal{M}^{\mathbb{T}}_{i,j})=\tilde{\mathcal{M}}^{\mathbb{T}}_{i,j}$. Thus, the set $\mathcal{J}_{i}$ of components of $\tilde{\mathcal{M}}_{i}^{\mathbb{T}}$ is in bijection with $\{i\}\sqcup\mathcal{J}^{\prime}_{i}$. In particular, every $\operatorname{\mathcal{E}}\in\tilde{\mathcal{M}}_{i,j}^{\mathbb{T}}$ is stable with resepct to the weight $w_{i}$.
@@
Recall the notation of Definition 1.5. Let $\operatorname{\mathcal{E}}_{0}\in\mathcal{M}_{i,j}^{\mathbb{T}}$ and write $\operatorname{\mathcal{E}}_{1}=\operatorname{\mathcal{E}}_{0}(S_{i})$. The proof is divided into three parts. In the first, we analyze the effect of a single Hecke transform on $\operatorname{\mathcal{E}}_{i}$. In the second part, we describe $\Psi_{i}(\operatorname{\mathcal{E}}_{1})$, and in the third and final part, we describe the Higgs bundle given by descent $\varXi_{i}(\operatorname{\mathcal{E}}_{0})=\Psi_{i}(\operatorname{\mathcal{E}}_{1})/\langle f\rangle.$
First part.
[…]
$s_{2}^{*}(b(s_{1}))(z)=\phi(z)dz.$ (55)
The Hecke transform is given as follows
$\operatorname{\mathcal{E}}_{1}^{\prime}=(L_{1}\oplus L_{2}(-x),\Phi_{b^{\prime}}),$ (56)
[…]
$\Phi_{b}^{\prime}=\begin{pmatrix}0&0\\ z^{-1}\phi(z)&0\end{pmatrix}\otimes dz,\quad(\epsilon_{j,x}=+1).$ (57)
The case that $\epsilon_{j,x}=-1$ is similar, except that in this case the modified bundle takes the form $\operatorname{\mathcal{E}}=L_{1}(-x)\oplus L_{2}$ and the modified Higgs field is of the following form (with similar notation)
$\Phi_{b}^{\prime}=\begin{pmatrix}0&0\\ z\phi(z)&0\end{pmatrix}\otimes dz,\quad(\epsilon_{j,x}=-1).$ (58)
We see that in each case $\Phi_{b}^{\prime}$ can be described as the off-diagonal Higgs field associated with the section $b^{\prime}=\iota^{-1}\circ b\circ\iota$, and we may write $\Phi_{b}^{\prime}=\Phi_{b^{\prime}}$. Moreover, in each case, it is clear that if we let $(F_{ab})_{1\leq a,b\leq 2}\subset\mathcal{O}_{X,x}$ be the matrix of regular functions that describes the lift $f_{\operatorname{\mathcal{E}}_{1}}$ with respect to the local frame $(s_{1},s_{2})$, then we see that $F_{12}$ and $F_{21}$ are constantly equal to zero. It follows that $(F_{ab})_{1\leq a,b\leq 2}$ is diagonal, and from remark 3, we deduce that the matrix of $F_{E^{\prime}}$ with respect to the frame $(s_{1}^{\prime},s_{2}^{\prime})$ is also diagonal. It follows that the new frame $(s_{1}^{\prime},s_{2}^{\prime})$ may be used in a potential next Hecke tranformation, and this will be used below.
Second part.
We now iterate the above analysis to describe $\Psi_{i}(\operatorname{\mathcal{E}}_{1})$. Define holomorphic line bundles defined as follows
$\Psi_{i,1}(L_{1}):=L_{1}\left(-\sum_{x\in X^{f}:\epsilon_{j,x}=-}b_{i}(x)\cdot x\right),$ (59)
$\Psi_{i,2}(L_{2}):=L_{2}\left(-\sum_{x\in X^{f}:\epsilon_{j,x}=+}b_{i}(x)\cdot x\right).$ (60)
We see by induction that
$\Psi_{i}(\operatorname{\mathcal{E}}_{1})=\prod_{x\in\operatorname{\tilde{\mathcal{D}}}_{i}}\Psi_{x}^{b_{i}(x)}(\operatorname{\mathcal{E}}_{1})=(\Psi_{i,1}(L_{1})\oplus\Psi_{i,2}(L_{2}),\Phi_{\Psi_{i}(b)}),$ (61)
where the first equality is simply the definition of $\Psi_{i}$, and where $\Phi_{\Psi_{i}(b)}$ is the off-diagonal Higgs field associated with the meromorphic section
$\Psi_{i}(b)\in H^{0}(X,\Psi_{i,1}(L_{1})^{-1}\otimes\Psi_{i,2}(L_{2})\otimes K_{X}),$ (62)
which is conjugate to $b$ on $X\setminus\operatorname{\tilde{\mathcal{D}}}_{i}$, and which near each $x\in\operatorname{\tilde{\mathcal{D}}}_{i}$ is of the following form
$(s^{\prime\prime}_{2})^{*}(\Psi_{i}(b)(s_{1}^{\prime\prime}))(z)=z^{-\epsilon_{j,x}b_{i}(x)}\phi(z)dz,$ (63)
where $(s^{\prime\prime}_{2},s^{\prime\prime}_{1})$ is a pair of holomorphic frames of the holomorphic lines bundles defining the underlying holomorphic vector bundle of the Higgs bundle and $z$ is a local coordinate centered at $x$.
Third part.
We now describe the effect of descent. By construction, the holomorphic line bundles $\Psi_{i,a}(L_{a}),a\in\{1,2\},$ and the Higgs field $\Phi_{\Psi(b)}$ are amenable to descent, and we write
$L_{+}:=\Psi_{i,2}(L_{2})/\langle f\rangle,$ (64)
$L_{-}:=\Psi_{i,1}(L_{1})/\langle f\rangle,$ (65)
so that $\varXi_{i}(\operatorname{\mathcal{E}}_{0})=(L_{+}\oplus L_{-},\Phi_{\Psi_{i}(b)}/\langle f\rangle)$. We now argue that
$\Psi_{i}(b)/\langle f\rangle\in H^{0}(\tilde{X},L_{+}L_{-}^{-1}K_{\tilde{X}}(\operatorname{\tilde{\mathcal{D}}}_{i}-\operatorname{\tilde{\mathcal{D}}}_{ij})).$ (66)
Recall the $\langle f\rangle$-invariant divisor $R_{i}=\mathcal{D}_{i}+(1-p)(X^{f}\setminus\mathcal{D}_{i}).$ By equation (19) we have that
$\Psi_{i}(b)\in H^{0}(X,\Psi_{i,1}(L_{1})^{-1}\Psi_{i,2}(L_{2})K_{X}(R_{i}))^{\langle f\rangle},$ (67)
and by construction, and as argued in the proof of Theorem LABEL:thm:M^f, we have that
$\Psi_{i}(b)/\langle f\rangle\in H^{0}(\tilde{X},L_{+}L_{-}^{-1}K_{\tilde{X}}(\operatorname{\tilde{\mathcal{D}}}_{i})).$ (68)
To show (66) we will show that $\Psi_{i}(b)$ belong to the image of the natural morphism (54). We will now analyze the zero divisor of $\Psi_{i}(b)$. Recall that by Lemma 1.2 there exists an effective divisor $\tilde{d}$ on $\tilde{X}$ of degree $l_{j}$ such that the zero divisor of $b$ is of the form
$(b)=\pi^{*}(\tilde{d})+j.$ (69)
It follows from (63) that the zero divisor of $\Psi_{i}(b)$ takes the form
$\begin{split}(\Psi_{i}(b))&=(b)+\sum_{x\in X^{f}}(1+\epsilon_{j,x}b_{i}(x))\cdot x-(p-1)(X^{f}/\mathcal{D}_{i})\\ &=\pi^{*}(\tilde{d})+j+\sum_{x\in X^{f}}(1+\epsilon_{j,x}b_{i}(x))\cdot x-(p-1)(X^{f}/\mathcal{D}_{i}).\end{split}$ (70)
Thus, to show that $\Psi_{i}(b)$ descends to a holomorphic section of (66) as desired, we must show that
$\pi^{*}(\operatorname{\tilde{\mathcal{D}}}_{ij})=j+\sum_{x\in\mathcal{D}_{i}}(1-\epsilon_{j,x}b_{i}(x))\cdot x-(p-1)(X^{f}\setminus\mathcal{D}_{i}).$ (71)
For each $x\in X^{f}$ denote by $d_{ij}(x)\in\mathbb{Z}$ the coefficient of $x$ in the divisor defined by the right hand side of (71). We must show that for each $x\in X^{f}$ we have that
$d_{ij}(x)=\begin{cases}p,&\text{if $\epsilon_{j,x}=-1$,}\\ 0,&\text{otherwise}.\end{cases}$ (72)
For $x\in X^{f}\setminus\mathcal{D}_{i}$, we have $i_{x}=0$, and therefore it follows directly from (39) that $j_{x}=p-1$. Hence $d_{ij}(x)=0$ in this case. Assume now that $x\in\mathcal{D}_{i}$ with $\epsilon_{j,x}=1$. Then $i_{x}\not=0\mod p$. From (40) and the fact that $n_{x}m_{x}=-1$, we obtain $2i_{x}m_{x}=1+j_{x}\mod p$. As $j\in\{0,...,p-1\}$ we observe that $1+j_{x}\in\{1,...,p\}$. However, as $2i_{x}m_{x}=1+j_{x}\mod p$ and as $2i_{x}m_{x}\not=0\mod p$, we have $1+j_{x}\in\{1,...,p-1\}$. Therefore $2i_{x}m_{x}]_{p}=1+j_{x}$. As $b_{i}(x)=[2i_{x}m_{x}]_{p}$ by definition, it follows that
$d_{ij}(x)=j_{x}+(1-\epsilon_{j,x}b_{i}(x))=j_{x}+(1-(1+j_{x}))=0,\quad(\epsilon_{j,x}=1).$ (73)
Assume now that $\epsilon_{j,x}=-1$. Then (40) and the fact that $n_{x}m_{x}=-1$ implies that $2i_{x}m_{x}=-(1+j_{x})\mod p$. As $j_{x}\in\{0,...,p-1\}$, this implies that $b_{i}(x)=[2i_{x}m_{x}]_{p}=p-(1+j_{x})$. Hence
$d_{ij}(x)=j_{x}+(1-\epsilon_{j,x}b_{i}(x))=j_{x}+(1+(p-(1+j_{x})))=p,\quad(\epsilon_{j,x}=-1).$ (74)
This finishes the argument that equation (71) holds.
@@
Finally, the fact that every $\operatorname{\mathcal{E}}\in\tilde{\mathcal{M}}_{i,j}^{\mathbb{T}}$ is stable with respect to $w_{i}$ follows from Theorem LABEL:thm:M^f, the fact that $\mathcal{M}^{\mathbb{T}}_{i,j}$ is a sub-variety of $\mathcal{M}_{i}$ and the identity $\varXi(\mathcal{M}^{\mathbb{T}}_{i,j})=\tilde{\mathcal{M}}^{\mathbb{T}}_{i,j}$. ∎
@@
From Corollary 1.1 it follows that every connected component is either a smooth moduli space of stable parabolic vector bundles on $\tilde{X}$ of rank two, or a principal-$\operatorname{Jac}(\tilde{X})[2]$ bundle over a symmetric power of $\tilde{X}$. Consider first the case of moduli spaces of stable bundles. The intersection pairing of these moduli spaces have been thoroughly studied following Witten’s work [76]. In the so-called coprime case, formulae for all pairings were proven in rank two in [70] and in all ranks in [49]. Subsequently, a version of these formulae for a class of parabolic moduli spaces including $\tilde{\mathcal{N}}_{i}$ were proven in [59]. Consider the case of the intersection pairings on the cohomology ring of symmetric powers of a Riemann surface. These rings are generated by MacDonald classes [58] and all the relations in the cohomology ring was completely described in [58]. This was used in [70] to compute certain characteristic numbers.